12. For each active ingredient, the amount required to make 30 capsules is Diphenhydramine HCl 750 mg Phenyltoloxamine citrate 900 mg Acetaminophen 9.75 g

Weight of 1 capsule full of Lactose = 475 mg Weight of 1 capsule full of Diphenhydramine HCl = 400 mg Weight of 1 capsule full of Phenyltoloxamine citrate = 375 mg Weight of 1 capsule full of Acetaminophen = 425 mg Determining the Amount of Lactose Required for making prescription for one capsule:

PLEASE MAY I HAVE ANSWER WITH BRIEF EXPLANATION
Answer #1

12. Determination of Amount of Lactose required for making the prescription for one capsule:

Step 1:

As per prescription,

Diphenhydramine HCl 25mg

Phenyltoloxamine citrate 30mg

Acetaminophen 325mg

Step 2: Calculate the percent of each ingredient

Then find the percent of each of the capsule filled by the active ingredient

Weight of 1capsule of Diphenhydramine HCl = 400mg.

Then percent of Diphenhydramine HCl = 25/400 X 100 = 6.25%

Weight of 1 capsule of Phenyltoloxamine Citrate = 375mg.

Then percent of Phenyltoloxamine Citrate = 30/375 X 100 = 8%

Weight of 1 capsule of Acetaminophen = 325mg

Then percent of Acetaminophen = 325 / 425 X 100 = 76.74%.

By adding all these three ingredient percent = (6.25 +8+ 76.74) = 90.72%

Step 3: Calculate the lactose percent

So the remaining capsule 9.28% ( 100 - 90.72) will be filled by the lactose in one capsule.

Calculate the amount of lactose for one capsule

Amount of Lactose = 9.28/100 X 475 = 44.08

The amount of Lactose required for making prescription of one capsule is 44mg.